Find two numbers whose sum is 10 and product is 18.
Answer by mananth(16946)

let the numbers be x & y
x+y= 10
xy=18
x=18/y
...
plug value of x in first equation
18/y + y =10
(18+y^2)/y =10
18+y^2=10y
y^2-10y+18=0
..
find y1 y2 the root os the equation using quadratic formuls
..
discriminant =28
..
y1 =((-(-10)+sqrt(28))/2
y1= 7.64
...
y2=((-(-10)-sqrt(28))/2
y2=2.36
..
the numbers are 7.64 & 2.36
...
m.ananth@hotmail.ca